OpenStudy (anonymous):

3(x-5)=5x+9 show steps

OpenStudy (anonymous):

\[3x-15=5x+9\] \[-15-9=5x-3x\] \[-24=2x\] \[x=\frac{-24}{2}=-12\]

OpenStudy (anonymous):

\[3x-15=5x+9\] \[3x-5x=9+15\] \[-2x=24\] \[{x=24/-2}\] \[x=-12\] remember a(b+c) = ab+ac

OpenStudy (anonymous):

photo shot1790 have u got the question or still any problem